# WaveDiffusion: Exploring Full Waveform Inversion via Joint Diffusion in the Latent Space

### v2

@@ 3.2 Evaluating Generated Samples with FID
Figure 5: t-SNE visualization of features for both real (blue) and generated (orange) data. Results of seismic data (col 1) and velocity map (col 2) for two joint diffusion models.
@@
Figure 6 visualizes results from the model with two VQs, while samples from the model with one VQ are shown in Appendix A.2. In the top row of each figure, we select samples structurally similar to FlatVel-B (FVB), where seismic inputs exhibit perfect symmetry, and the generated velocity maps maintain this symmetry, demonstrating the model’s ability to respect geometric constraints. Beyond reproducing individual OpenFWI subsets, we also observe cases where features from multiple datasets are fused, as seen in the ninth row of Figure 6, showcasing the diversity and fidelity of the generated seismic-velocity pairs. These results confirm that the joint diffusion model generalizes well across different data distributions, effectively capturing structural coherence and producing reliable outputs for seismic data-velocity map generation.
3.3 Diffusion Improves FWI Performance
@@
For this experiment, the Stage 1 model functions as an image-to-image translation network, analogous to BigFWI-B. In contrast, the diffusion model in Stage 2 refines the latent representation using the last 10 backward denoising steps, aiming to improve reconstruction accuracy. The refinement process is performed as follows: at each denoising step, we randomly sample 100 new latent vectors $z_{t}$ and decode them into seismic data $\hat{s}$. Each decoded $\hat{s}$ is compared with the corresponding input seismic data $s$ by computing the L2 distance. The sample with the lowest L2 distance is selected as the current $z_{t}$ for the next backward denoising step. This iterative refinement continues until reaching $z_{0}$, at which point the final reconstructed seismic-velocity pair is obtained.
Improvement by diffusion: Our results, as shown in Figure 1 for metrics comparison, demonstrate that the Stage 1 model achieves inversion performance comparable to BigFWI-B across all datasets. Additionally, applying the diffusion model to refine the latent space leads to slight but consistent improvements in reconstruction accuracy. Visualization of examples can be found in Figure 7. These findings suggest that the latent representations learned by the Stage 1 model already encode meaningful physical structures, and the diffusion process further enhances them by guiding reconstructions toward physically valid solutions.
Figure 7: Inversion performance visualization. Inverted examples from CVB (row 1), FFB (row 2), and FVB (row 3) subsets. (Col 1) ground truth; (col 2) 2VQ without diffusion; (col 3) 2VQ with diffusion; (col 4) BigFWI-B.
For a more in-depth analysis of numerical comparisons, including RMSE, MAE, and SSIM metrics, we refer readers to Appendix A.3.
3.4 Training Data Augmentation
Table 2: Quantitative results of InversionNet performance (MAE) training on pure generated data. Performance of InversionNet trained on jointly generated and original OpenFWI data.
@@ 3.4 Training Data Augmentation
Table 3: Quantitative results of InversionNet performance (MAE) for partial real data and partial real + generated data. Performance of InversionNet trained on partial OpenFWI data and partial real data plus generated data.
@@
We evaluate the effectiveness of WaveDiffusion-generated samples in augmenting training data, assessing how well they supplement the original OpenFWI dataset.
Real data vs. generated data: To examine the standalone quality of generated data, we train InversionNet (Wu & Lin, 2019) exclusively on WaveDiffusion-generated samples, ensuring the dataset matches the scale of the OpenFWI training set. The model is then tested on the OpenFWI test set, and results are compared against the state-of-the-art BigFWI-B (Jin et al., 2024) baseline. As shown in Table 2, models trained purely on generated data underperform compared to those trained on real data, indicating a gap in the fidelity of the generated samples.
Partial real data vs. partial real + gen: We further investigate the impact of mixing generated data and partial real data for training, focusing on samples generated by the model with one shared VQ. Specifically, we evaluate two settings: Gen+$n\%$Real, where generated data is combined with a small portion ($10\%$ or $1\%$) of real data, and $n\%$Real, where models are trained exclusively on the same small subset of real data.
Table 3 presents the results for both settings. Incorporating even a small fraction of real data (Gen+$n\%$Real) significantly improves performance. Gen+10%Real demonstrates a substantial improvement over using 10%Real alone, with metrics approaching the BifFWI-B. Similarly, Gen+1%Real outperforms 10%Real on most datasets, particularly on more complex data. These findings highlight the effectiveness of augmenting generated data with even a small amount of real data. In contrast, the $n\%$Real models exhibit the weakest performance, emphasizing the limitations of relying solely on a small dataset for effective training.
A prediction visualization is shown in Figure 8. These results demonstrate that while generated samples can effectively supplement small datasets, the inclusion of even a small portion of real data is crucial for achieving optimal results. More prediction visualizations are shown in Appendix A.4.
@@ 4 Related works
Traditional physics-based FWI: Traditional FWI methods aim to reconstruct subsurface velocity models by iteratively minimizing the difference between observed and simulated seismic data, typically using gradient-based optimization methods. The key challenge lies in solving the wave equation, which governs wave propagation through the Earth. While effective, these methods are computationally expensive and sensitive to factors such as the quality of the initial velocity model, noise in the data, and cycle-skipping issues—where the inversion algorithm converges to incorrect solutions due to poor starting models or insufficient low-frequency data (Tarantola, 1984; Virieux & Operto, 2009). Techniques such as adaptive waveform inversion (Warner & Guasch, 2016) and multiscale FWI (Bunks et al., 1995) have been developed to reduce the risk of cycle-skipping and improve convergence by progressively introducing higher-frequency data. These techniques frame FWI as a conditional generation problem, relying on physical equations as the computational foundation (Virieux et al., 2017; Warner & Guasch, 2016).
Data-driven approaches to FWI: In recent years, machine learning approaches have been increasingly explored for FWI. Convolutional Neural Networks (CNNs) have shown promise in learning image-to-image mappings from seismic data to velocity models, bypassing the need for iterative solvers. Encoder-decoder architectures, such as those used in InversionNet (Wu & Lin, 2019) and VelocityGAN (Zhang et al., 2019), have demonstrated the ability to predict velocity maps from seismic data while reducing computational costs by learning implicit relationships between the two modalities. Richardson’s work (Richardson, 2018) further illustrated that deep learning models could predict velocity models efficiently. However, these approaches are still treating the FWI problem as an image-to-image translation task or a conditional generation problem (Zhu et al., 2019; Wang et al., 2023a). Recent work on neural operators (Li et al., 2020, 2023) offers a more flexible approach by learning operators that map between the two modalities in a revered direction, i.e., predicting seismic data given velocity maps. While neural operator methods demonstrate a strong capacity in modality mapping, their reliance on direct image-to-image translation prevents them from capturing the intrinsic physical constraints embedded in the latent space.
@@
Generative models in FWI: Generative models, particularly Generative Adversarial Networks (GANs) and their variants, have emerged as alternatives to traditional CNN-based methods for FWI. These models aim to learn the latent representations of seismic data and velocity models, enabling the generation of synthetic training data or even direct inversion (Goodfellow et al., 2020). Vector Quantized GANs (VQGANs) (Esser et al., 2021), in particular, have been explored for their ability to generate high-quality modalities, such as images, audios, videos, etc. Such models can be tuned for imaging one physical modality (e.g., velocity) given another (e.g., seismic) (Zhang et al., 2019).
@@
Recent work has focused on Latent Diffusion Models (LDMs) (Ho et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2022), which refine latent space representations through a diffusion process. LDMs iteratively denoise latent variables, progressively improving the quality of generated samples. While these models can produce realistic-looking data, they often generate new samples of one single modality at a time. Thus, it is difficult for them to generate multiple modalities using one generative model as they lack the physical consistency to the governing PDEs that describe the relationship between these modalities. Diffusion models have been applied to FWI by Wang et al. (Wang et al., 2023a), who used them to generate prior distributions for plausible velocity models as a regularization term. Their method still treats seismic data and velocity maps separately, limiting its ability to generate physically consistent seismic-velocity pairs.
@@ 5 Conclusion
In this work, we revisit FWI from a generative perspective and propose WaveDiffusion, a diffusion-based framework that refines the latent space constructed by encoder-decoder models. We demonstrate that standard image-to-image translation FWI models do not ensure physical consistency in their learned latent representations. To address this, we employ a joint diffusion model in the shared latent space that learns to score deviations from the governing PDE, progressively transforming arbitrary latent points into physically consistent solutions, ensuring that generated seismic-velocity pairs naturally satisfy the governing PDE. Additionally, our experiments show that the diffusion model can improve the encoder-decoder performance of FWI tasks, and the generated data can serve as the augmentation of the training set for data-driven FWI models, particularly in scenarios with limited real data availability. Our findings provide a new perspective on bridging generative modeling with physics-based problem-solving, paving the way for diffusion models to enhance scientific discovery and computational physics applications.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Network Details and Training Hyperparameters
A.1.1 encoder-decoder Model
The encoder and decoder use convolutional encoder-decoder branches that are constructed by ResNet blocks (He et al., 2016) for seismic and velocity data. The channel multipliers of the ResNet blocks are set to [1, 2, 2, 4, 4] for velocity maps and [1, 2, 2, 4, 4, 4, 4, 8, 8] for seismic data. The resolution for velocity maps is 64, while for seismic data, it is [1024, 64]. The size of the latent feature map $z$ is [16, 16], and the number of residual blocks is set to 3. The model was trained with a base learning rate of $4.5\times 10^{-4}$. It uses an embedding dimension of 32 and an embedding codebook size of 8192. It employs a perceptual loss combined with a discriminator. The discriminator starts training at step 50001 with a discriminator weight of 0.5 and a perceptual weight of 0.5.
@@ A.1.3 Training Hyperparameters
We trained our models on 128 NVIDIA GH200 GPUs. Training required approximately 8000 GPU hours for the first-stage encoder-decoder model and 12000 GPU hours for the joint diffusion model.
@@ A.2 Joint Generation Examples
A.3 Statistical Comparison of Our Models and BigFWI-B
We present the detailed statistical comparison between the BigFWI-B model and our first-stage encoder-decoder model and the corresponding diffusion models on the inversion tasks for the velocity maps across all the datasets of OpenFWI.
A.4 InversionNet Results Visulization

### v3

@@ 3.2 Evaluating Generated Samples with FID
Figure 7 visualizes results from the model with two VQs, while samples from the model with one VQ are shown in Appendix A.2. In the top row of each figure, we select samples structurally similar to FlatVel-B (FVB), where seismic exhibits perfect symmetry, and the velocity maps maintain this symmetry, demonstrating the model’s ability to respect geometric constraints. Beyond reproducing individual OpenFWI subsets, we also observe cases where features from multiple datasets are fused, as seen in the ninth row, showcasing the diversity of the generated results. These results confirm that the joint diffusion model generalizes well across different data distributions, effectively capturing structural coherence and producing reliable outputs for seismic-velocity data generation.
To illustrate the potential of our approach beyond acoustic FWI, we also conducted experiments on the elastic FWI dataset, $\mathbf{\mathbb{E}^{FWI}}$ Feng et al. (2023). The results are reported in Appendix A.4.
3.3 Diffusion Improves FWI
@@
For this experiment, the Stage 1 model functions as an image-to-image translation network, analogous to BigFWI-B. In contrast, the diffusion model in Stage 2 refines the Stage 1 model’s latent representation using the last 10 backward denoising steps, aiming to improve reconstruction accuracy. The refinement process is performed as follows: at each denoising step, we randomly sample 100 new latent vectors $z_{t}$ and decode them into seismic data $\hat{s}$. Each decoded $\hat{s}$ is compared with the corresponding input seismic data $s$ by computing the L2 distance. The sample with the lowest L2 distance is selected as the current $z_{t}$ for the next backward denoising step. This iterative refinement continues until reaching $z_{0}$, at which point the final reconstructed seismic-velocity pair is obtained.
Figure 8: Inversion performance visualization. Inverted examples from CVB (row 1), FFB (row 2), and FVB (row 3) subsets.
Improvement by diffusion: Our results, as shown in Figure 2, demonstrate that the Stage 1 model achieves inversion performance comparable to BigFWI-B across all datasets. Additionally, applying the diffusion model to refine the latent space leads to slight but consistent improvements. Visualization of examples can be found in Figure 8. These findings suggest that the latent representations learned in Stage 1 already encode meaningful physical structures, and the diffusion process further enhances them by guiding reconstructions toward physically valid solutions. For a more in-depth analysis of numerical results, we refer readers to Appendix A.5 and A.6.
3.4 Evaluating the Utility of Generated Data for FWI
To evaluate the effectiveness of generated samples, we conduct a series of downstream experiments using image-to-image FWI models. Specifically, we assess how well the generated data can supplement the original OpenFWI dataset in low-data regimes and as an additional source for fine-tuning.
Real data vs. generated data: We first assess the standalone quality of the generated data by training InversionNet Wu and Lin (2019) exclusively on samples generated by WaveDiffusion, using a training set that matches the scale of the original OpenFWI dataset. The model is then evaluated on the OpenFWI test set, and the results are compared with the state-of-the-art BigFWI-B Jin et al. (2024) baseline. As shown in Table 2, models trained purely on generated samples underperform compared to those trained on real data, indicating a gap in fidelity despite structural consistency.
[…]
Figure 9 illustrates prediction visualizations, while additional results are provided in Appendix A.7. Overall, these results demonstrate that while generated samples can effectively supplement small datasets, the inclusion of even a small portion of real data is crucial for achieving optimal results.
Providing additional data with challenging samples: In addition to low-data scenarios, we evaluate the benefit of using generated samples as additional training data to fine-tune a pretrained model. Specifically, we select the 65K most challenging generated samples, on which BigFWI-B performs worst, and partition them into 55K for training and 10K for testing. We then fine-tune BigFWI-B for 20 additional epochs on the combined dataset (original OpenFWI + generated challenging samples). For a fair comparison, we compare with a model fine-tuned only on the original OpenFWI training set for the same number of epochs. As shown in Table 4, fine-tuning with generated challenging samples yields consistent improvements across all ten datasets in OpenFWI, along with a significant boost in performance on the generated test samples. This suggests that WaveDiffusion not only produces physically meaningful samples but also exposes hard modes that enhance model performance.
Table 2: MAE of InversionNet training on pure generated data.
@@ 3.4 Evaluating the Utility of Generated Data for FWI
Table 3: MAE of InversionNet for partial real data and partial real + generated data.
@@
Table 4: MAE of InversionNet for fine-tuning on the combined dataset versus OpenFWI-only. LDM denotes performance on 1K generated challenging samples.
Dataset FVA FVB CVA CVB FFA FFB CFA CFB SA SB LDM
Combined data 0.0062 0.0220 0.0338 0.0927 0.0097 0.0716 0.0160 0.1256 0.0512 0.0548 0.1900
OpenFWI only 0.0064 0.0220 0.0339 0.0933 0.0102 0.0718 0.0163 0.1259 0.0513 0.0549 0.2120
@@ 4 Related Works
Data-driven approaches to FWI: Recent learning-based methods bypass iterative solvers by training neural networks to map seismic data to velocity models directly. Encoder-decoder architectures such as InversionNet Wu and Lin (2019) and VelocityGAN Zhang et al. (2019) reduce computational cost by learning implicit mappings between the two modalities. With higher efficiency than physics-based FWI, these methods treat FWI as an image-to-image translation task Richardson (2018); Zhu et al. (2019). More recently, neural operators Li et al. (2020; 2023) provide a flexible alternative for predicting seismic data from velocity models, with strong capacity for modality transformation. In Appendix A.9, we review traditional physics-driven FWI methods.
@@
Generative models in FWI: Generative models, particularly Generative Adversarial Networks (GANs) and their variants, have emerged as alternatives to traditional CNN-based methods for FWI. These models aim to learn the latent representations of seismic data and velocity models, enabling the generation of synthetic training data or even direct inversion Goodfellow et al. (2020). Vector Quantized GANs (VQGANs) Esser et al. (2021), in particular, have been explored for their ability to generate high-quality modalities, such as images, audios, videos, etc. Such models can be tuned for imaging one physical modality (e.g., velocity) given another (e.g., seismic) Zhang et al. (2019).
@@
Diffusion models have also been applied to FWI Wang et al. (2023a), who used them to generate prior distributions for plausible velocity models and incorporated with a regularization process through a Plug-and-Play (PnP)-style framework. However, their approach treats seismic data and velocity maps separately and relies on explicit forward modeling. Recent work on Latent Diffusion Models (LDMs) Ho et al. (2020); Dhariwal and Nichol (2021); Rombach et al. (2022) refines latent representations through a diffusion process. While effective for generating realistic single-modality data, these models are difficult to produce multiple physical consistency modalities jointly.
5 Limitation
While this work introduces a novel generative perspective on seismic and velocity data modeling, several limitations remain. First, we provide no formal guarantees or theoretical characterization of PDE satisfaction. Our contribution is to highlight and empirically quantify the emergent property and tendency of improved physical adherence in the generated seismic-velocity pairs from the joint latent diffusion. Second, despite the diffusion process promoting physical consistency, the PDE satisfaction is not perfect. This is a common limitation of data-driven methods that approximate complex physical systems. Finally, although we conducted preliminary experiments on the elastic FWI setting, our study remains primarily focused on the acoustic case. A more systematic investigation of more wave physics or other inverse problems remains an important direction for future research.
@@ 6 Conclusion
In this work, we revisit the relationship of seismic and velocity data from a generative perspective and propose WaveDiffusion, a diffusion-based framework for physics-consistent latent modeling. We demonstrate that standard image-to-image translation FWI models do not ensure physical consistency in their learned latent representations. To address this, we employ a joint diffusion model in the shared latent space that learns to evaluate deviations from the governing PDE, progressively transforming arbitrary latent points into physically consistent solutions, ensuring that generated seismic-velocity pairs naturally satisfy the governing PDE. Additionally, our experiments show that the diffusion model can improve the performance of FWI tasks, and the generated data can serve as a supplement to the training set for data-driven FWI models. Our findings provide a new perspective on bridging generative modeling with physics-based problem-solving, paving the way for diffusion models to enhance scientific discovery and computational physics applications.
@@ A.1 Network Details and Training Hyperparameters
A.1.1 Encoder-Decoder Model with Shared Latent Space and Vector Quantization
Shared Latent Space Construction: In the first stage of the WaveDiffusion framework, we build a shared latent space using a VQ-VAE architecture that comprises one encoder and two decoders. The encoder $E$ maps the input seismic data $s$ into a continuous latent representation $z_{\text{cont}}$, which is then discretized by a vector quantization (VQ) layer into $z$. Two decoders, $D_{s}$ and $D_{v}$, are jointly trained to reconstruct the seismic data ($\hat{s}$) and the corresponding velocity maps ($\hat{v}$) from the same latent code $z$. This joint training ensures that the latent space captures the intrinsic dependencies between the two modalities, thereby enabling paired generation.
To clarify the training procedure, Algorithm 1 outlines our approach.
Algorithm 1 Training the Shared Latent Space in WaveDiffusion
0: Training pairs $\{(s,v)\}$
1:Initialize: Encoder $E$, decoders $D_{s}$ and $D_{v}$, vector quantizer (VQ), and discriminator (for perceptual loss).
2:for each training iteration do
3:  Compute continuous latent code: $z_{\text{cont}}=E(s)$.
4:  Quantize latent code: $z=\text{VQ}(z_{\text{cont}})$.
5:  Reconstruct seismic data: $\hat{s}=D_{s}(z)$.
6:  Reconstruct velocity map: $\hat{v}=D_{v}(z)$.
7:  Compute reconstruction losses:
$\mathcal{L}_{s}=\|s-\hat{s}\|_{1},\quad\mathcal{L}_{v}=\|v-\hat{v}\|_{1}.$
8:  Compute perceptual loss $\mathcal{L}_{\text{perc}}$ (using a discriminator or LPIPS loss).
9:  Form the total loss:
$\mathcal{L}=\mathcal{L}_{s}+\mathcal{L}_{v}+\mathcal{L}_{\text{perc}}.$
10:  Backpropagate $\mathcal{L}$ and update $E$, $D_{s}$, $D_{v}$, and the VQ layer.
11:end for
Vector Quantization: The VQ layer discretizes the continuous latent vectors into a finite codebook (with an embedding dimension of 32 and a codebook size of 8192). This discretization helps capture structured patterns and long-range dependencies between seismic data and velocity maps.
Architectural Details: The encoder and decoders are implemented as convolutional encoder-decoder networks using ResNet blocks He et al. (2016). For velocity maps, the channel multipliers are set to $[1,2,2,4,4]$ with an output resolution of 64. For seismic data, the multipliers are $[1,2,2,4,4,4,4,8,8]$ with resolutions $[1024,64]$. The latent feature map $z$ has dimensions $[16,16]$, and 3 residual blocks are used. The model is trained with a base learning rate of $4.5\times 10^{-4}$ and employs a perceptual loss combined with an adversarial loss. The discriminator begins training at step 50001 with weights of 0.5 for both the discriminator and perceptual loss components.
@@ A.2 Joint Generation Examples
A.3 Evaluating Latent Distribution Sampling: Random Sampling and Interpolation
We conducted two experiments to address the concern that poor reconstructions result merely from latent distribution mismatch.
  1. 1.
    Random Sampling Experiment: We first normalized the training latent vectors (300K samples) to a Gaussian (zero mean, std=1) and then sampled 5000 new points from this distribution. The decoded velocity maps from these new points show merged features from the training data. However, the corresponding reconstructed seismic data (s^)($\hat${s}) exhibit significantly larger differences from the ground truth seismic data $s$ (obtained via finite-difference simulation on the decoded velocity maps). This confirms that poor reconstructions are not merely due to sampling from a mismatched distribution. Visualizations can be found at Figure 11.
  2. 2.
    Interpolation Experiment: We interpolated between two training latent points (A and B, 5000 pairs in total) using five different weights (i.e., $0.1A+0.9B$, $0.3A+0.7B$, $0.5A+0.5B$, $0.7A+0.3B$, and $0.9A+0.1B$). While the decoded velocity maps smoothly merge features of A and B, the resulting seismic data differences $(\hat{s}-s)$ remain large. This indicates that the auto-encoder’s latent space does not inherently support paired generation, underscoring the necessity of the diffusion process to refine the shared latent space for PDE-consistent outputs. Visualizations can be found at Figure 12.
To clarify, the two experiments above do not contain any diffusion process, and are only performed using the first stage encoder-decoder network. The results are summarized in Table 5:
These results demonstrate that encoder-decoder sampling (both random and interpolation) yields relatively high MSE values, which indicates most latent space points do not correspond to valid PDE solutions. The diffusion process significantly reduces the MSE, confirming that the diffusion process guides latent codes toward outputs that satisfy the PDE much more effectively.
Figure 11: Random Sampling Experiment.
Figure 12: Interpolation Experiment.
Table 5: Evaluating latent distribution sampling, comparing diffusion with random Sampling and interpolation.
Methods Seismic Averaged MSE
Random Sampling 0.0082
Interpolation Sampling 0.0065
Diffusion Denoising 0.0038
A.4 Extension to Elastic FWI
We further evaluated the generality of our framework by conducting experiments on one sub-dataset (FFA) from the $\mathbf{\mathbb{E}^{FWI}}$ dataset Feng et al. (2023). To adapt our network to the elastic setting, we stacked the two seismic components ($x$ and $z$) along the channel dimension, and likewise stacked the $P$- and $S$-wave velocity maps. Using the trained model, we generated 6K samples and report the FID scores across all four modalities in Table 6.
FID (VP) FID (VS) FID (Data_x) FID (Data_z)
Value 195.91 301.77 3.40 1.15
Table 6: FID scores for elastic FWI across different modalities.
These results are comparable to those obtained in the acoustic FWI experiments, suggesting that our framework extends naturally to the elastic case. In terms of PDE satisfaction, we observed a similar denoising trend: the average $L_{2}$ deviation for the $x$-component of seismic data decreased from $7.98\times 10^{-4}$ to $7.80\times 10^{-4}$ during denoising, while for the $z$-component it decreased from $1.8\times 10^{-5}$ to $0.8\times 10^{-5}$. Notably, these deviations are even lower than in the acoustic setting, which we attribute to training on a single sub-dataset, where more consistent data patterns make learning easier.
A.5 Statistical Comparison of Our Models and BigFWI-B
We present the detailed statistical comparison in Table 9 of the BigFWI-B model and our first-stage encoder-decoder models and the corresponding diffusion models on the inversion tasks for the velocity maps across all the datasets of OpenFWI.
A.6 Ablation of Latent Refinement via Diffusion model
In addition to our initial experiment, we conducted a further analysis as follows: We randomly sampled 1,000 latent vectors from the Gaussian distribution centered around the latent position corresponding to the target seismic data (from the Style-B dataset) with a slight oscillation ($5\%$) from the target seismic data encoded latent point. Next, for each sample, we decoded the seismic data (s^)($\hat${s}) and computed the MSE loss with respect to the target seismic data $s$. We then selected the latent point with the lowest MSE loss and evaluated the paired velocity map using RMSE, MAE, and SSIM metrics.
The results are summarized in Table 7. An example comparison can be found at Figure 13. These results indicate that the encoder-decoder random sampling cannot improve over the baseline, and the diffusion model FWI consistently achieves lower MAE and RMSE, as well as a higher SSIM, demonstrating an improvement by the diffusion process. The random sampling around the baseline latent point by the encoder-decoder network only even decreases the performance. This confirms that the diffusion process effectively refines the latent representation, yielding samples that better satisfy the governing PDE and thus improve FWI performance.
Table 7: Ablation of Latent Refinement via Diffusion model, comparing with ocillation sampling around baseline on the SB dataset.
Methods MAE RMSE SSIM
Starting Baseline 0.0574 0.0882 0.9026
Oscillation Sampling around Baseline 0.1199 0.1582 0.7722
Diffusion Model FWI 0.0423 0.0873 0.9069
Figure 13: Visualization of Latent Refinement via Diffusion model.
A.7 InversionNet Results Visulization